Hello, how do I I find the equation of the line joining (-3,-3) and (4,8)? I need to leave the answer in general form. I am struggling to figure this out, so any help is greatly appreciated! :)
Answer by rfer(16322)

slope m=(8--3)/(4--3)
m=11/7
---
y=mx+b
8=11/7*4+b
8=44/7+b
8-44/7=b
56/7-44/7=b
12/7=b
----------------
1 4/7x+y=1 5/7